# From Age-Structured Trophic Networks to Applied Control : Stabilization and Harvesting Strategies for Non-Transitive Competition and the Dynamics of Mosquitoes

### v1

@@ 3.1.1.2. Stability for $\psi_{i}\neq 0$
(3.86) $V_{G}(\eta,\psi)=V_{3}(\eta_{2},z_{1},z_{2})\;+\;\frac{\gamma_{1}}{\sigma_{1}}h(G_{1}(\psi_{1}))\;+\;\frac{\gamma_{2}}{\sigma_{2}}h(G_{2}(\psi_{2}))\;+\;\frac{\gamma_{3}}{\sigma_{3}}h(G_{3}(\psi_{3}))$
@@
(3.88) $\mathcal{K}_{3}=\left\{(\eta,\psi)\in\mathbb{R}^{3}\times\mathcal{S}^{3}\ \middle|\ \begin{aligned} &\eta_{1}\leq\ln\left(\frac{\gamma_{1}}{C_{1}\lambda_{1}}\right),\qquad&&\gamma_{1}>C_{1}\lambda_{1},\\[4.0pt] &\eta_{2}\leq\ln\left(\frac{\gamma_{2}}{C_{2}\lambda_{2}}\right),\qquad&&\gamma_{2}>C_{2}\lambda_{2},\\[4.0pt] &\eta_{3}\leq\ln\left(\frac{\gamma_{3}}{C_{3}\lambda_{3}}\right),\qquad&&\gamma_{3}>C_{3}\lambda_{3},\\[6.0pt] &u^{*}+c_{3o}\phi_{1}(z_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z_{1})+(c_{2}-1-\theta)\phi_{3}(z_{1})\\ &+\frac{c_{2}}{\lambda_{1}}\phi_{3}(z_{1})\phi_{1}(z_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-(c_{1}+1)\phi_{2}(\eta_{2})\\ &+\frac{\theta\phi_{3}(z_{1})}{\phi_{1}(z_{2})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z_{1})\right)>0.\end{aligned}\right\}.$
[…]
@@
    (3.89) $V_{3}(\eta_{2},z_{1},z_{2})=\theta\omega_{2}(\eta_{2},z_{1})+\lambda_{1}(e^{z_{2}}-1-z_{2})$
@@
(3.91) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-c_{30}\phi^{2}_{1}(z_{2})$
$+\phi_{3}(z_{1})\left(\theta c_{2}\,\phi_{3}(z_{1})+\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\theta\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\theta(c_{1}-1)\phi_{2}(\eta_{2})-\theta\phi_{3}(z_{1})\right)+\phi_{2}(\eta_{2})\left(\theta c_{1}\,\phi_{2}(\eta_{2})+(c_{1}+1)\phi_{1}(z_{2})-\phi_{1}(z_{2})\right)$
$+\phi_{1}(z_{2})\phi_{3}(z_{1})\left(\theta\frac{c_{2}}{\lambda_{1}}\phi_{3}(z_{1})-(c_{2}-1-\theta)-\frac{c_{2}}{\lambda_{1}}\phi_{1}(z_{2})+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\right)+\left(\theta\phi_{3}(z_{1})-\phi_{1}(z_{2})-\theta\phi_{2}(\eta_{2})\right)\phi_{3}(\eta_{3})+\left(\phi_{1}(z_{2})-\theta\phi_{3}(z_{1})\right)\phi_{1}(\eta_{1})$
$+\underbrace{\left(\theta\phi_{3}(z_{1})-\phi_{1}(z_{2})-\theta\phi_{2}(\eta_{2})\right)}_{A_{3}}|\hat{\phi}_{3}-\phi_{3}(\eta_{3})|+\underbrace{\left(\phi_{1}(z_{2})-\theta\phi_{3}(z_{1})\right)}_{A_{1}}|\hat{\phi}_{1}-\phi_{1}(\eta_{1})|\underbrace{-\phi_{1}(z_{2})}_{A_{2}}|\hat{\phi}_{2}-\phi_{2}(\eta_{2})|$
@@
(3.92) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-c_{3}\phi^{2}_{1}(z_{2})+A_{3}\left(\hat{\phi}_{3}-\phi_{3}(\eta_{3})\right)+A_{1}\left(\hat{\phi}_{1}-\phi_{1}(\eta_{1})\right)+A_{2}\left(\hat{\phi}_{2}-\phi_{2}(\eta_{2})\right)+\mathcal{R}$
@@
(3.93) $\mathcal{R}=\phi_{3}(z_{1})\left(\theta c_{2}\,\phi_{3}(z_{1})+\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\theta\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})-\theta(c_{1}-1)\phi_{2}(\eta_{2})-\theta\phi_{3}(z_{1})\right)+\phi_{2}(\eta_{2})\left(\theta c_{1}\,\phi_{2}(\eta_{2})+c_{1}\phi_{1}(z_{2})\right)$
$+\phi_{1}(z_{2})\phi_{3}(z_{1})\left(\theta\frac{c_{2}}{\lambda_{1}}\phi_{3}(z_{1})-(c_{2}-1-\theta)-\frac{c_{2}}{\lambda_{1}}\phi_{1}(z_{2})+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\right)+\left(\theta\phi_{3}(z_{1})-\phi_{1}(z_{2})-\theta\phi_{2}(\eta_{2})\right)\phi_{3}(\eta_{3})+\left(\phi_{1}(z_{2})-\theta\phi_{3}(z_{1})\right)\phi_{1}(\eta_{1})$
@@
(3.94) $\begin{cases}\phi_{1}(\eta_{1})=c_{2}\phi_{3}(z_{1})+\phi_{1}(z_{2})+\frac{c_{2}}{\lambda_{1}}\phi_{1}(z_{2})\phi_{3}(z_{1}),\\ \phi_{3}(\eta_{3})=c_{1}\phi_{2}(\eta_{2})+\phi_{3}(z_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})\end{cases}$
@@
(3.96) $\dot{V}_{3}(\eta_{2},z_{1},z_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-c_{3}\phi^{2}_{1}(z_{2})+A_{3}\left(\hat{\phi}_{3}-\phi_{3}(\eta_{3})\right)+A_{1}\left(\hat{\phi}_{1}-\phi_{1}(\eta_{1})\right)+A_{2}\left(\hat{\phi}_{2}-\phi_{2}(\eta_{2})\right)$
@@
(3.97) $\dot{V}_{G}(\eta,\psi)\leq\dot{V}_{3}(\eta_{2},z_{1},z_{2})\;-\;\gamma_{1}(e^{G_{1}}-1)\;-\;\gamma_{2}(e^{G_{2}}-1)\;-\;\gamma_{3}(e^{G_{3}}-1)$
(3.98) $\dot{V}_{G}(\eta,\psi)\leq-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-c_{3}\phi^{2}_{1}(z_{2})+|A_{1}||\hat{\phi}_{1}-\phi_{1}|+|A_{2}||\hat{\phi}_{2}-\phi_{2}|+|A_{3}||\hat{\phi}_{3}-\phi_{3}|-\;\gamma_{1}(e^{G_{1}}-1)-\gamma_{2}(e^{G_{2}}-1)-\gamma_{3}(e^{G_{3}}-1).$
@@
(3.99) $\dot{V}_{G}(\eta,\psi)\leq-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-c_{3}\phi^{2}_{1}(z_{2})+|A_{1}||\phi_{1}(\eta_{1})+\lambda_{1}|(e^{v_{1}}-1)+|A_{2}||\phi_{2}(\eta_{2})+\lambda_{2}|(e^{v_{2}}-1)+|A_{3}||\phi_{3}(\eta_{3})+\lambda_{3}|(e^{v_{3}}-1)$
@@
(3.100) $\dot{V}_{G}(\eta,\psi)\leq-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z_{1})-c_{3}\phi^{2}_{1}(z_{2})+\left(C_{1}|\phi_{1}(\eta_{1})+\lambda_{1}|-\gamma_{1}\right)(e^{G_{1}}-1)+\left(C_{2}|\phi_{2}(\eta_{2})+\lambda_{2}|-\gamma_{2}\right)(e^{G_{2}}-1)$
@@
(3.101) $\begin{cases}\eta_{1}\leq\underbrace{\ln\left(\frac{\gamma_{1}}{C_{1}\lambda_{1}}\right)}_{H_{1}},\\ \\ \eta_{2}\leq\underbrace{\ln\left(\frac{\gamma_{2}}{C_{2}\lambda_{2}}\right)}_{H_{2}},\\ \\ \eta_{3}\leq\underbrace{\ln\left(\frac{\gamma_{3}}{C_{3}\lambda_{3}}\right)}_{H_{3}},\end{cases}\qquad\text{with}\;\begin{cases}\gamma_{1}>C_{1}\lambda_{1},\\ \\ \gamma_{2}>C_{2}\lambda_{2},\\ \\ \gamma_{3}>C_{3}\lambda_{3}.\end{cases}$
we get global asymptotic stability.∎
Proposition 3.13.
There exists $c>0$ such that a connected component of the sublevel set
$\mathcal{L}_{c}:=\{(\eta,\psi)\mid V_{G}(\eta,\psi)\leq c\}$
is contained in $\mathcal{K}_{3}$.
[…]
    We select an initial vector
    $(\eta^{0},\psi^{0}),\qquad\text{with }\eta^{0}=(\varepsilon,0,0),\;\;\psi^{0}=(0,0,0),$
    where $\varepsilon\neq 0$ is a sufficiently small real number, chosen such that $0<\varepsilon<\min\Big{\{}\ln\tfrac{\gamma_{1}}{C_{1}\lambda_{1}},\;\ln\tfrac{\gamma_{2}}{C_{2}\lambda_{2}},\;\ln\tfrac{\gamma_{3}}{C_{3}\lambda_{3}}\Big{\}}.$ Since $\phi_{1}$ is continuous with $\phi_{1}(0)=0$, one can further choose $\varepsilon>0$ small enough to guarantee $|c_{3o}\phi_{1}(\varepsilon)|<\tfrac{u^{*}}{2}.$ Therefore,
    $u^{*}+c_{3o}\phi_{1}(\varepsilon)>\tfrac{u^{*}}{2}>0,$
    which shows that $(\eta^{0},\psi^{0})\in\mathcal{K}_{3}$. In particular, this implies $\mathcal{K}_{3}\neq\varnothing$. Next, by continuity we have
    $V_{G}(\eta^{0},\psi^{0})=V_{3}(0,0,\varepsilon)+\sum_{i=1}^{3}\frac{\gamma_{i}}{\sigma_{i}}\,h(G_{i}(\psi_{i}^{0})).$
    Since $V_{3}(0,0,\varepsilon)\rightarrow 0$ as $\varepsilon\rightarrow 0$, and $h(0)=0$ whenever $G_{i}(\psi_{i}^{0})=0$, it follows that $V_{G}(\eta^{0},\psi^{0})$ can be made arbitrarily small.
    Hence, for any given $c>0$, there exists $\varepsilon>0$ sufficiently small such that $V_{G}(\eta^{0},\psi^{0})\leq c.$ Consequently, $(\eta^{0},\psi^{0})\in\mathcal{L}_{c}\cap\mathcal{K}_{3},$ which proves that $\mathcal{L}_{c}\cap\mathcal{K}_{3}\neq\varnothing.$
@@
The aim of the constructed fictitious controls is to stabilize, step by step and recursively, subsystems of model until the global control $u$ is synthesized. Concretely, each fictitious control is designed to partially stabilize a given state or subsystem. For example, in transformation (3.43), the state $\eta_{2}$ is partially stabilized so as to serve as a reference for the stabilization of $\eta_{3}$. Then the signal $z_{1}$ partially stabilizes $\eta_{3}$, and the process proceeds stepwise until the general control $u$ is obtained. Each stage guarantees partial stability of the considered subsystem, which allows one to infer stabilization of the full system.

### v2

@@ 3.1.1.2. Stability for $\psi_{i}\neq 0$
(3.86) $V_{G}(\eta,\psi)=V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})\;+\;\frac{\gamma_{1}}{\sigma_{1}}h(G_{1}(\psi_{1}))\;+\;\frac{\gamma_{2}}{\sigma_{2}}h(G_{2}(\psi_{2}))\;+\;\frac{\gamma_{3}}{\sigma_{3}}h(G_{3}(\psi_{3}))$
@@
(3.88) $\mathcal{K}_{3}=\left\{(\eta,\psi)\in\mathbb{R}^{3}\times\mathcal{S}^{3}\ \middle|\ \begin{aligned} &\eta_{1}\leq\ln\left(\frac{\gamma_{1}}{C_{1}\lambda_{1}}\right),\qquad&&\gamma_{1}>C_{1}\lambda_{1},\\[4.0pt] &\eta_{2}\leq\ln\left(\frac{\gamma_{2}}{C_{2}\lambda_{2}}\right),\qquad&&\gamma_{2}>C_{2}\lambda_{2},\\[4.0pt] &\eta_{3}\leq\ln\left(\frac{\gamma_{3}}{C_{3}\lambda_{3}}\right),\qquad&&\gamma_{3}>C_{3}\lambda_{3},\\[6.0pt] &u^{*}+c_{3o}\phi_{1}(z^{3}_{2})-\theta\frac{c_{2}}{\lambda_{1}}\phi^{2}_{3}(z^{3}_{1})+(c_{2}-1-\theta)\phi_{3}(z^{3}_{1})\\ &+\frac{c_{2}}{\lambda_{1}}\phi_{3}(z^{3}_{1})\phi_{1}(z^{3}_{2})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-(c_{1}+1)\phi_{2}(\eta_{2})\\ &+\frac{\theta\phi_{3}(z^{3}_{1})}{\phi_{1}(z^{3}_{2})}\left(\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})+(c_{1}-1)\phi_{2}(\eta_{2})+\,\phi_{3}(z^{3}_{1})\right)>0.\end{aligned}\right\}.$
  • Proof of Theorem 3.13 :
@@
    (3.89) $V_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=\theta\omega_{2}(\eta_{2},z^{3}_{1})+\lambda_{1}(e^{z^{3}_{2}}-1-z^{3}_{2})$
@@
    (3.91) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})-c_{30}\phi^{2}_{1}(z^{3}_{2})$
    $+\phi_{3}(z^{3}_{1})\left(\theta c_{2}\,\phi_{3}(z^{3}_{1})+\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\theta\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-\theta(c_{1}-1)\phi_{2}(\eta_{2})-\theta\phi_{3}(z^{3}_{1})\right)+\phi_{2}(\eta_{2})\left(\theta c_{1}\,\phi_{2}(\eta_{2})+(c_{1}+1)\phi_{1}(z^{3}_{2})-\phi_{1}(z^{3}_{2})\right)$
    $+\phi_{1}(z^{3}_{2})\phi_{3}(z^{3}_{1})\left(\theta\frac{c_{2}}{\lambda_{1}}\phi_{3}(z^{3}_{1})-(c_{2}-1-\theta)-\frac{c_{2}}{\lambda_{1}}\phi_{1}(z^{3}_{2})+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\right)+\left(\theta\phi_{3}(z^{3}_{1})-\phi_{1}(z^{3}_{2})-\theta\phi_{2}(\eta_{2})\right)\phi_{3}(\eta_{3})+\left(\phi_{1}(z^{3}_{2})-\theta\phi_{3}(z^{3}_{1})\right)\phi_{1}(\eta_{1})$
    $+\underbrace{\left(\theta\phi_{3}(z^{3}_{1})-\phi_{1}(z^{3}_{2})-\theta\phi_{2}(\eta_{2})\right)}_{A_{3}}\left(\hat{\phi}_{3}-\phi_{3}(\eta_{3})\right)+\underbrace{\left(\phi_{1}(z^{3}_{2})-\theta\phi_{3}(z^{3}_{1})\right)}_{A_{1}}\left(\hat{\phi}_{1}-\phi_{1}(\eta_{1})\right)\underbrace{-\phi_{1}(z^{3}_{2})}_{A_{2}}\left(\hat{\phi}_{2}-\phi_{2}(\eta_{2})\right).$
@@
    (3.92) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})-c_{3}\phi^{2}_{1}(z^{3}_{2})+A_{3}\left(\hat{\phi}_{3}-\phi_{3}(\eta_{3})\right)+A_{1}\left(\hat{\phi}_{1}-\phi_{1}(\eta_{1})\right)+A_{2}\left(\hat{\phi}_{2}-\phi_{2}(\eta_{2})\right)+\mathcal{R},$
@@
    (3.93) $\mathcal{R}=\phi_{3}(z^{3}_{1})\left(\theta c_{2}\,\phi_{3}(z^{3}_{1})+\theta\frac{c_{1}}{\lambda_{3}}\phi^{2}_{2}(\eta_{2})-\theta\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2})-\theta(c_{1}-1)\phi_{2}(\eta_{2})-\theta\phi_{3}(z^{3}_{1})\right)+\phi_{2}(\eta_{2})\left(\theta c_{1}\,\phi_{2}(\eta_{2})+c_{1}\phi_{1}(z^{3}_{2})\right)$
    $+\phi_{1}(z^{3}_{2})\phi_{3}(z^{3}_{1})\left(\theta\frac{c_{2}}{\lambda_{1}}\phi_{3}(z^{3}_{1})-(c_{2}-1-\theta)-\frac{c_{2}}{\lambda_{1}}\phi_{1}(z^{3}_{2})+\frac{c_{1}}{\lambda_{3}}\phi_{2}(\eta_{2})\right)+\left(\theta\phi_{3}(z^{3}_{1})-\phi_{1}(z^{3}_{2})-\theta\phi_{2}(\eta_{2})\right)\phi_{3}(\eta_{3})+\left(\phi_{1}(z^{3}_{2})-\theta\phi_{3}(z^{3}_{1})\right)\phi_{1}(\eta_{1}).$
@@
    (3.94) $\begin{cases}\phi_{1}(\eta_{1})=c_{2}\phi_{3}(z^{3}_{1})+\phi_{1}(z^{3}_{2})+\frac{c_{2}}{\lambda_{1}}\phi_{1}(z^{3}_{2})\phi_{3}(z^{3}_{1}),\\ \phi_{3}(\eta_{3})=c_{1}\phi_{2}(\eta_{2})+\phi_{3}(z^{3}_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{3}(z^{3}_{1})\phi_{2}(\eta_{2}),\end{cases}$
@@
    (3.96) $\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})-c_{3}\phi^{2}_{1}(z^{3}_{2})+A_{3}\left(\hat{\phi}_{3}-\phi_{3}(\eta_{3})\right)+A_{1}\left(\hat{\phi}_{1}-\phi_{1}(\eta_{1})\right)+A_{2}\left(\hat{\phi}_{2}-\phi_{2}(\eta_{2})\right)$
@@
    (3.97) $\dot{V}_{G}(\eta,\psi)\leq\dot{V}_{3}(\eta_{2},z^{3}_{1},z^{3}_{2})\;-\;\gamma_{1}(e^{G_{1}}-1)\;-\;\gamma_{2}(e^{G_{2}}-1)\;-\;\gamma_{3}(e^{G_{3}}-1)$
    (3.98) $\dot{V}_{G}(\eta,\psi)\leq-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})-c_{3}\phi^{2}_{1}(z^{3}_{2})+|A_{1}||\hat{\phi}_{1}-\phi_{1}|+|A_{2}||\hat{\phi}_{2}-\phi_{2}|+|A_{3}||\hat{\phi}_{3}-\phi_{3}|-\;\gamma_{1}(e^{G_{1}}-1)-\gamma_{2}(e^{G_{2}}-1)-\gamma_{3}(e^{G_{3}}-1).$
@@
    (3.99) $\dot{V}_{G}(\eta,\psi)\leq-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})-c_{3}\phi^{2}_{1}(z^{3}_{2})+|A_{1}||\phi_{1}(\eta_{1})+\lambda_{1}|(e^{v_{1}}-1)+|A_{2}||\phi_{2}(\eta_{2})+\lambda_{2}|(e^{v_{2}}-1)+|A_{3}||\phi_{3}(\eta_{3})+\lambda_{3}|(e^{v_{3}}-1)$
@@
    (3.100) $\dot{V}_{G}(\eta,\psi)\leq-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta c_{2}\,\phi_{3}^{2}(z^{3}_{1})-c_{3}\phi^{2}_{1}(z^{3}_{2})+\left(C_{1}|\phi_{1}(\eta_{1})+\lambda_{1}|-\gamma_{1}\right)(e^{G_{1}}-1)+\left(C_{2}|\phi_{2}(\eta_{2})+\lambda_{2}|-\gamma_{2}\right)(e^{G_{2}}-1)$
@@
    (3.101) $\begin{cases}\eta_{1}\leq\ln\left(\frac{\gamma_{1}}{C_{1}\lambda_{1}}\right),\\ \\ \eta_{2}\leq\ln\left(\frac{\gamma_{2}}{C_{2}\lambda_{2}}\right),\\ \\ \eta_{3}\leq\ln\left(\frac{\gamma_{3}}{C_{3}\lambda_{3}}\right),\end{cases}\hskip 18.49988pt\text{with}\;\begin{cases}\gamma_{1}>C_{1}\lambda_{1},\\ \\ \gamma_{2}>C_{2}\lambda_{2},\\ \\ \gamma_{3}>C_{3}\lambda_{3}.\end{cases}$
    we get global asymptotic stability.
@@
The aim of the constructed fictitious controls is to stabilize, step by step and recursively, subsystems of model until the global control $u$ is synthesized. Concretely, each fictitious control is designed to partially stabilize a given state or subsystem. For instance, in transformation (3.42), the state $\eta_{2}$, serving as a reference for system stabilization, is partially stabilized by relation (3.48). Subsequently, $\eta_{3}$ is in turn partially stabilized by $\eta_{1}$ via the signal $z^{3}_{1}$ (see Condition (3.56)). The procedure runs inductively: each step guarantees partial stability of the corresponding subsystem, and ultimately yields an appropriate global control law $u$ from which stabilization of the closed-loop system follows.
@@
Regarding the mathematical formulation, we confirm that $z^{3}_{2}\neq 0$ by Lemma 3.23.